# Photonic-Electronic Integrated Circuits for High-Performance Computing and AI Accelerator

### v1

@@ II-C Toward Fully-functional EPALU
Besides the aforementioned advantages of high bandwidth, low latency, and reduced power consumption of optical logic units, unique multiplexing techniques play a pivotal role in further improving computing capacity and performance. An example is the WDM-based electronics-photonic arithmetic logic unit (EPALU) (Fig. 3.b) [25]. The arithmetic logic unit (ALU) performs arithmetic and bitwise operations and serves as an essential component in modern computing systems. Within ALUs, the adder plays a critical role, with its logic being expressed as follows:
@@
where $A$, $B$, and $C$ represent the two operands and carry respectively; $p$ and $g$ mean propagation and generation, and the subscripts indicate the specific bits. Based on this expression, the scalable electro-optic carry propagation adder (CPA) has been developed by Ying et al. [58], with the schematic shown in the Fig. 3.a. For a $N$-bit CPA, the carry output from one stage is the carry input to the next stage, thus the final result cannot be calculated until the carry has rippled through all stages. As an optimized architecture, the carry select adder (CSA) splits $N$-bit operands into $n$ $m$-bit CPA ($N=m\times n$). It speeds up addition by computing two possible outcomes for each $m$-bit CPA simultaneously—assuming carry-in values of ”0” and ”1”—and then selects the correct result based on the actual carry-in using multiplexers (MUXs) [70]. As a trade-off, CSAs require two sets of circuits with different carry inputs. However, in the optical domain, this can be efficiently implemented with a single set of PIC, where different carry signals are encoded into two wavelengths using WDM (Fig. 3.c). Beyond arithmetic addition, this multifunctional architecture enables the execution of addition, subtraction, comparison, and bitwise operations with various input combinations. Additionally, the E-O shifter within the EPALU architecture has also been experimentally demonstrated (Fig. 3.d) [71].
From the perspective of computer architecture, the ALU operates under the directives issued by the control unit, with its inputs being fetched from memory. Upon completing a designated operation, the ALU’s output is then stored back in memory to be accessed for subsequent computations. While the latency of arithmetic units has been significantly reduced by optical computing units, the speed of data access and control signal transmission may introduce new bottlenecks for the entire computing system. Therefore, the exploration of high-speed interconnects between different modules within an optical computing system is of significant importance as well [72, 73]. In Ref [73], a decoder and multiplexer designed for the EPALU architecture have been demonstrated, achieving data transportation and processing both within and between modules at a speed of 20 Gb/s (Fig. 3.e).
Figure 3: EPALU architecture for high-performance digital computing. (a) Schematic of a 2-bit CPA using E-O logic [58]. (b) Schematic of EPALU with their building blocks and data path. (c) The architecture of the WDM-based $N$-bit multifunctional processing unit consists of a $(p,g)$ generation unit (PGU), n sets of $m$-bit optical carry propagation networks (OCPNs) and an array of photodetectors (PDs) along with a network of electronic multiplexer units (MUXU) and an electronic sum generation unit (SGU) [25]. With different input combinations, EPALU can perform various logic functions (right). (d) Layout of a 2-bit barrel shifter using microdisk add-drop switch array [71]. (e) Schematic of an $n$-$2^{n}$ electro-photonic decoder and $2^{n}$-$n$ photonic-electronic multiplexer, where $s_{i}$ refers to the electrical input signal.
@@ III PIC-Based Analog computing for AI: Fundamentals and Implementation
Undoubtedly, modern AI, functioning on digital computing systems, has achieved significant progress in diverse fields and has even exceeded human performance in specific tasks. With the advancements in integrated photonics, the PIC platform emerges as a compelling candidate for AI accelerators. Leveraging the optical logic gates and computing units detailed in Section II, some studies have illustrated the capability of PICs to perform tensor operations, including accumulation, dot products, and matrix-vector multiplications (MVMs), for diverse machine learning applications within the digital domain [74, 75, 76]. These studies have highlighted the inherent advantages of optical computing in terms of latency, speed, and power efficiency. However, within the photonic-electronic platform, the digital representation can encounter challenges stemming from hardware complexity overhead and speed reduction caused by the sampling and digitization into binary streams processed by logic units. These challenges are especially significant in the context of high-throughput or high-precision tensor operations for various machine-learning tasks.
On the other hand, the human brain, operating as an analog signal ”processor”, is estimated to perform at a rate of $10^{18}$ multiply-accumulate (MAC)/s with a power consumption estimated at just $\sim$10-20 W [77, 78], demonstrating remarkable efficiency compared to the substantial energy requirements of cutting-edge AI. This efficiency potentially be attributed to the parallel processing capability and reduced precision requirements inherent in analog computing systems. While the mechanisms of brain function remain incompletely understood, there is growing interest in incorporating analog computing into machine learning. Before discussing the details of implementing optical analog computing for AI, it is helpful to provide a concise overview of artificial neural networks (ANNs) and the neuron model. The schematic of an artificial neuron with a basic synaptic model is illustrated in Fig. 4, where x, y, and w represent the inputs from the pre-synaptic neuron, post-synaptic output, and weights of the connection, respectively. Activation functions $\sigma(\cdot)$, such as sigmoid, ReLU, and the leaky integrate-and-fire (LIF) function for SNNs, introduce nonlinearity into the model along with various engineering considerations [79, 80].
@@ III-A Programmable modulation for optical analog computing
In analog AI accelerators, both inputs x and weights w could correspond to a higher level of programming resolution, in contrast to the binary values in digital computing circuits. The hardware implementation of the above process using PIC requires the reconfigurable programming of network parameters, which relies on the modulation of optical components. While the weights in a trained model may remain fixed during the inference process, it is still necessary to calibrate the network through modulation due to fabrication variations of PIC. A number of modulation mechanisms have been developed, among which tuning the effective refractive index $n_{\text{eff}}$ of waveguides is a widely adopted approach in ONNs.
@@ III-A2 Field-effect tuning mechanism
Except for the thermo-optic effect, the $n_{\text{eff}}$ can also be tuned by electric fields. In silicon photonics, a straightforward approach involves doping the silicon waveguide and applying an external electric field to manipulate the carrier concentration and tune $n_{\text{eff}}$. A variety of modulators on the PIC platform operating in carrier-injection (forward-bias p-i-n), carrier-depletion (reverse-bias p-n junction), and carrier-accumulation (MOSCAP) mode have been widely demonstrated (Fig. 5.b) [83, 84, 85, 86, 87, 88, 89]. These CMOS-compatible mechanisms allow gigahertz-level tuning speeds, making them suitable for high-speed input encoding in ONNs. Specifically, in depletion mode, the bandwidth is determined by the majority of carriers’ dynamics, which are not limited by the slower processes of carrier generation and recombination [86]. Furthermore, while the depletion-mode modulator remains a volatile device, it exhibits low static power consumption attributed to the reverse-biased junction. Compared to thermo-optic devices, the free-carrier effect-based modulators typically require longer modulation length and a larger footprint, primarily due to the lower efficiency (as evaluated by voltage-length product $V_{\pi}L$) or the necessity for traveling-wave electrodes to achieve high-speed modulation. Expanding beyond the conventional silicon platform, an alternative approach is utilizing materials with significant electro-optic effects—such as the free-carrier plasma dispersion, Pockels effect, Kerr effect, and Quantum Confined Stark Effect (QCSE)—for the core or cladding of waveguides. The modulation efficiency of these devices is intrinsically linked to the material properties. Prominent examples include III–V semiconductors [90, 88, 89, 91], lithium niobate [92], and some polymers [93, 94]. A comprehensive discussion of all these mechanisms exceeds the scope of this review. Interested readers are referred to relevant reviews and reports for further information [95, 96, 97].
@@ III-A3 Non-volatile modulation
Each mechanism discussed above is a volatile process and requires continuous power supply even for infrequent or static programming tasks. Non-volatile phase change materials (PCMs) offer a unique opportunity for these scenarios. PCMs have two switchable phases, i.e. amorphous and crystalline states, with drastically different $n_{\text{eff}}$, and can achieve reversible phase transitions within various temperature ranges. For the PIC platform, heating can be achieved by an integrated pulse microheater or alternatively, through light signals themselves, enabling ’all-optical’ modulation. Chen et al. developed several non-volatile silicon photonics modulators with Sb2S3 PCM cladding and experimentally demonstrated 5-bit multilevel programming with a high cyclability $>$1600 switching events (Fig. 5.c) [98]. Besides, all-optical SNNs using chalcogenide PCM GeSbTe (GST) also have been reported [34], and more details will be discussed in Section. III-C2. A brief summary of modulation techniques with Si waveguide is shown in Table I.
Figure 5: Modulation techniques of photonic-electronic devices on PICs. (a) Optical micrograph and the normalized transmission curve of a thermo-optic MZI. Each arm of the MZI incorporates a microheater as a phase shifter [81]. While both arms can be used to modulate the output signal, typically, one is used to set the modulator output at the quadrature point for a high extinction ratio in monotonic modulation. (b) Schematics and micrographs of free-carrier effect-based modulators working on carrier-injection mode [99], carrier-depletion mode [87], and MOSCAP-driven tuning mechanism [91]. (c) Top, optical micrograph of an MRR modulator loaded with 10 $\mu$m long 20-nm-thick Sb2S3 and doped silicon PIN heater. Bottom, normalized transmission spectra of MRR when switching between two phases of Sb2S3 [98].
@@
$f_{3\text{dB}}$
Data Rate
Energy Efficiency
Active Length
Modulator Ref
Thermo-optics $\sim$0.023 $<$1
$\sim$4 KHz
$\sim$1.36 mW*
$\sim$170 $\mu$m
MZM [81]
Carrier Injection 0.036 12
16 GHz
10 Gb/s
51 mW* / 5 pJ/bit
200 $\mu$m
MZM [84]
Carrier Depletion 1.62–2.05 3.9
30 GHz
44 Gb/s
2.84 pJ/bit
1000 $\mu$m
MZM [87]
Si-MOSCAP $\sim$0.8 9
50 GHz
$>$100 Gb/s
R = 15 $\mu$m
MRR [88]
III-V-MOSCAP 0.09 1
2.2 GHz
32 Gb/s
250 $\mu$m
MZM [91]
Graphene-MOSCAP 0.28 $\sim$7
5 GHz
10 Gb/s
1 pJ/bit
300 $\mu$m
MZM [100]
@@ 10 Gb/s
III-B Implementations of Photonic Tensor Core
As illustrated in Fig. 4, the interconnections in ANNs can be conceptualized as weights akin to synaptic coupling coefficients in biological systems. This analogy extends to representing these connections through tensor operations, thereby abstracting the complex interactions in a computationally manageable form. Building upon the aforementioned modulation mechanisms, various active devices and encoding mechanisms have been effectively utilized in photonic neurons. This section categorizes the ONNs from multiple perspectives, which aim to provide a comprehensive analysis and comparison of the implementations of photonic tensor core (PTC) from diverse optical components to the architectural design.
@@ III-B1 Coherent vs incoherent ONN
From the viewpoint of signal properties, ONNs can be classified into coherent and incoherent systems. Within coherent ONNs, both weights and inputs can be encoded in the complex plane, allowing multiplication through lossless interference. For instance, a pair of beam splitters and phase shifters in the form of MZIs are widely adopted for conducting linear operations in coherent ONNs  [101] due to the programmable amplitude and phase response. The schematic of 2$\times$2 MZI is shown in Fig. 6.a. Assuming that each beam splitter is an ideal 50:50 directional coupler (i.e., with a coupling coefficient $\kappa=1/\sqrt{2}$), the transition matrix of the MZI can be represented as shown in Eq.  (3).
@@
Here, $\phi_{1}$ and $\phi_{2}$ represent the phase changes as signals traverse two arms. In some cases, $\phi_{1}$ and $\phi_{2}$ are denoted as $\Delta\phi=\phi_{1}-\phi_{2}$ and 0, respectively, given that the response of MZI is predominantly governed by the phase difference between its two arms. If only one input port receives a signal $\bm{E}_{\textbf{in}}$, the field amplitudes and power intensities at the output port can be expressed as follows:
@@
This implies that the output signal’s phase and amplitude can be modulated by varying $\Delta\phi$ from 0 to $\pi$. The smooth transmission curve, stemming from the cosine term, enables high-resolution analog computing and enhances noise resistance. In a notable case, the two arms are encoded by opposite phase shifts, represented as $\phi\pm\Delta\phi/2$ for $\phi_{1}$ and $\phi_{2}$, respectively. When incorporated into Eq. (3), the Mach-Zehnder modulator (MZM) working on the “push-pull” mode allows for the independent modulation of field intensity while maintaining a constant phase.
@@
$U(2)$ $=\left[\begin{array}[]{cc}e^{-j\theta}&0\\[4.0pt] 0&1\\ \end{array}\right]\cdot\frac{1}{2}\left[\begin{array}[]{cc}e^{-j\Delta\phi}-1&j(e^{-j\Delta\phi}+1)\\[6.0pt] j(e^{-j\Delta\phi}+1)&-(e^{-j\Delta\phi}-1)\\ \end{array}\right]$ (10)
@@
Unitary matrices of rank $N$ can be decomposed into sets of $U(2)$ blocks, which can be implemented using cascaded 2$\times$2 MZIs to form a mesh structure. Despite the unitary nature of the transition matrix in an MZI-mesh, an arbitrary weight matrix for an ONN can be implemented by singular vector decomposition (SVD). The SVD approach was thoroughly discussed by Miller et al. in 2003 and first experimentally implemented in ONNs by Shen et al. as shown in Fig. 6.b [102, 29]. Specifically, an arbitrary real-valued matrix $\bm{M}$ can be decomposed into the form $\bm{U\Sigma V^{\dagger}}$, where $\bm{\Sigma}$ are a diagonal matrix and the remaining two are unitary matrices, which can be implemented with a set of tunable attenuators and MZI mesh, respectively.
Additionally, another characteristic of the devices based on phase modulation is their broad spectral bandwidth. During the modulation process, the phase change $\psi$ can be expressed as follows:
@@
Thanks to the phase encoding mechanism, coherent ONNs can perform multiplication with full-range input operands. Besides the SVD approach, a more intuitive example is that the sign and absolute value of the scalar can be encoded by the phase shifter and MZM separately. Following this methodology, Ref [104] presents a photonic neuron architecture enabling full-range dot product operations. However, this architecture also faces limitations in terms of complexity and scalability. Specifically, implementing SVD required a matrix pre-processing step for phase mapping, which consumes extra time and power. In addition, the number of MZI needed to implement escalates quadratically with matrix size, leading to scalability issues from a larger footprint and accumulated losses in cascaded blocks. Additionally, the phase control of the coherence network presents a challenge as well. Within the MZI-mesh architecture, each $U(2)$ operation demands a minimum of two active phase modulators. Without a reference signal, phase calibration becomes more complex than amplitude calibration, which can be straightforwardly measured using a photodetector, especially considering the fabrication deviation across individual devices and waveguides. However, some studies demonstrate the feasibility of addressing the phase calibration issue through on-chip training [105, 106].
Incoherent PICs offer an alternative approach that can be configured to execute tensor operations. In the absence of coherence requirements, incoherent ONN enables broader flexibility in employing multiplexing techniques. Because of its wavelength-dependent transmission characteristics, MRRs are widely employed for encoding weights and inputs in WDM-based incoherent ONNs. From the perspective of physical design, the compactness of MRR (achieving a radius <10 µm) affords a denser layout and enhanced scalability of ONN. As shown in Fig. 6.c, add-drop MRR is a four-port optical device that consists of a microring evanescently coupled to two bus waveguides. The transfer characteristics at the drop port and through port can be expressed as Eq. (15) and Eq. (16).

### v2

@@ II-C Toward a Fully-functional EPALU
Besides the aforementioned advantages of optical logic units, unique multiplexing techniques play a pivotal role in further improving computing capacity and performance. An example is the WDM-based electronics-photonic arithmetic logic unit (EPALU) (Fig. 3.a) [24]. The arithmetic logic unit (ALU), which performs arithmetic and bitwise operations, is an essential component of modern computing systems. Within ALUs, the full adder plays a critical role, with its logic being expressed as follows:
@@
where $A$, $B$, and $C$ represent the two operands and carry, respectively; $p$ and $g$ denote propagation and generation, and the subscripts indicate the specific bits. Based on these expressions, the scalable electro-optic carry propagation adder (CPA) has been developed [61], with the schematic shown in Fig. 3.c. For a $N$-bit CPA, the carry output from one stage is the carry input to the next stage, thus the final result cannot be calculated until the carry has rippled through all stages. As an optimized architecture, the carry select adder (CSA) splits $N$-bit operands into $n$ $m$-bit CPA ($N=m\times n$). It speeds up addition by computing two possible outcomes for each $m$-bit CPA simultaneously—assuming carry-in values of ‘0’ and ‘1’—and then selects the correct result based on the actual carry-in using multiplexers (MUXs) [75]. As a trade-off, CSAs require two sets of circuits with different carry inputs. However, in the optical domain, this can be efficiently implemented with a single optical path in the PIC, where different carry signals are encoded into two distinct wavelengths using WDM (Fig. 3.e). Beyond arithmetic addition, Ying et al. developed that the multifunctional architecture could perform addition, subtraction, comparison, and bitwise operations operating at 20 GB/s with various input combinations [24]. Based on time-space multiplexing, Zhang et al. demonstrated a photonic-electronic digital multiplier capable of processing up to 32 $\times$ 4-bit binary inputs at 25 Mbit/s [76]. Additionally, the E-O shifter within the EPALU architecture has also been experimentally demonstrated (Fig. 3.d) [77].
From the perspective of computer architecture, the ALU operates under the directives issued by electronic control units, with its inputs being fetched from memory. Upon completing a designated operation, the ALU’s output is then stored back in memory to be accessed for subsequent computations (Fig. 3.a). Although optical computing has significantly reduced the latency of arithmetic operations, data access and E-O/O-E conversions can create major bandwidth bottlenecks and serve as a significant source of energy consumption in the computing system. Therefore, the exploration of high-speed interconnects between different modules within an electronic–photonic microprocessor is important as well [78, 79]. In Ref [79], a decoder and multiplexer designed for the EPALU architecture have been demonstrated, achieving data transportation and processing at a speed of 20 Gb/s (Fig. 3.b). Another potential approach to reduce the time and energy costs of E-O/O-E conversion is through the implementation of various forms of optical memory, as discussed in Ref [80, 81].
[…]
@@ III PIC-Based Analog computing for AI: Fundamentals and Implementation
Undoubtedly, modern AI, functioning on digital computing systems, has achieved significant progress in diverse fields and has even exceeded human performance in specific tasks. With the advancements in integrated photonics, the PIC platform emerges as a compelling candidate for AI accelerators. Leveraging the optical logic gates and computing units detailed in Section II, some studies have illustrated the capability of PICs to perform tensor operations, including accumulation, dot products, and matrix-vector multiplications (MVMs), for diverse machine learning applications within the digital domain [82, 83, 84]. These studies have highlighted the inherent advantages of optical computing in terms of latency, speed, and power efficiency. However, the digital representation can encounter challenges stemming from hardware complexity overhead and speed reduction caused by the sampling and digitization into binary streams processed by logic units. These challenges are especially significant in the context of high-throughput or high-precision tensor operations for various machine-learning tasks.
On the other hand, the human brain, operating as an analog signal “processor”, is estimated to perform at a rate of $10^{18}$ multiply-accumulate (MAC)/s with a power consumption estimated at just $\sim$10-20 W [85, 86], demonstrating remarkable efficiency compared to the substantial energy requirements of cutting-edge AI. This efficiency can potentially be attributed to the parallel processing capability and reduced precision requirements inherent in analog computing. While the mechanisms of brain function remain incompletely understood, there is growing interest in incorporating analog computing into machine learning. Before discussing the details of implementing optical analog computing for AI, it is helpful to provide a concise overview of artificial neural networks (ANNs) and the neuron model. The schematic of an artificial neuron with a basic synaptic model is illustrated in Fig. 4, where x, y, and w represent the inputs from the pre-synaptic neuron, post-synaptic output, and weights of the connection, respectively. The activation functions $\sigma(\cdot)$, such as sigmoid, ReLU, and the leaky integrate-and-fire (LIF) function for SNNs, introduce nonlinearity into the model along with various engineering considerations [87, 88].
@@ III-A Programmable modulation for optical analog computing
In analog AI accelerators, both inputs x and weights w could correspond to a higher resolution, in contrast to the binary values in digital computing circuits. The hardware implementation of the above process using PICs requires the reconfigurable programming of network parameters, which relies on the modulation of optical components. While the weights in a trained model may remain fixed during the inference process, it is still necessary to calibrate the network through modulation due to the fabrication variations of PICs. A number of modulation mechanisms have been developed, among which tuning the effective refractive index $n_{\text{eff}}$ of waveguides is a widely adopted approach in ONNs.
@@ III-A2 Field-effect tuning mechanism
Except for the thermo-optic effect, the $n_{\text{eff}}$ can also be tuned by electric fields. In silicon photonics, a straightforward approach involves doping the silicon waveguide and applying an external electric field to manipulate the carrier concentration and tune $n_{\text{eff}}$. A variety of modulators operating in carrier-injection (forward-bias p-i-n), carrier-depletion (reverse-bias p-n junction), and carrier-accumulation (metal-oxide-semiconductor capacitor, MOSCAP) mode have been widely demonstrated (Fig. 5.b) [91, 92, 93, 94, 95, 96]. These CMOS-compatible mechanisms allow gigahertz-level tuning speeds, making them suitable for high-speed encoding in ONNs. Particularly, in depletion mode, the bandwidth is determined by the majority carriers’ dynamics, which are not limited by the slower processes of carrier generation and recombination [94]. Moreover, while the depletion-mode modulator remains a volatile device, it exhibits low static power consumption attributed to the reverse-biased junction. Compared to thermo-optic devices, the free-carrier effect-based modulators typically require longer modulation lengths and larger footprints, primarily due to their lower tunning efficiencies (as evaluated by voltage-length product $V_{\pi}L$) or the necessity for traveling-wave electrodes for high-speed modulation. Expanding beyond the conventional silicon platform, an alternative approach is the utilization of materials with significant electro-optic effects—such as the free-carrier plasma dispersions, the Pockels effect, the Kerr effect, and the Quantum Confined Stark Effect (QCSE)—for the core or cladding of waveguides. The modulation efficiency of these devices is intrinsically linked to the material properties. Prominent examples include III–V semiconductors [97, 96, 98], lithium niobate [99], and some polymers [100, 101]. A comprehensive discussion of all these mechanisms exceeds the scope of this review. Interested readers are referred to relevant reviews and reports for further information [102, 103, 104].
Most modulators discussed above have nonlinear transmission curves, primarily due to the E-O modulation mechanisms involved. Although the Pockels effect provides a linear change in $\Delta n$ in response to electric field intensity, the inherent transfer function of the modulator itself can still exhibit a nonlinearity (such as the sinusoidal response of MZI). To reduce undesired nonlinearity in computing architectures, several solutions have been proposed: (1) Introduce a nonlinear mapping during D/A conversion, although this may necessitate additional data processing; (2) Utilize the nearly linear segment of the transmission curve, at the tradeoff of a reduced modulation dynamic range; (3) Design PICs that compensate for nonlinearity by novel devices or modulation mechanisms [105, 106, 100, 107, 108]. Additionally, the impact of nonlinearity on system noise also needs consideration. Within modulators exhibiting strong nonlinearity and steep slopes, such as high-Q MRRs, slight noise or offsets in the drive signal can potentially result in significant deviations.
@@ III-A3 Non-volatile modulation
Each aforementioned mechanism is a volatile process and requires continuous power supply even for infrequent or static programming tasks. Non-volatile phase change materials (PCMs) offer a unique opportunity to avoid these scenarios. PCMs have two switchable phases, i.e. amorphous and crystalline states, with drastically different $n_{\text{eff}}$, and can achieve reversible phase transitions within various temperature ranges. For the PIC platform, heating can be achieved by an integrated pulse microheater or alternatively, through light signals themselves, enabling ’all-optical’ modulation. Chen et al. developed several non-volatile silicon photonics modulators with Sb2S3 PCM cladding and experimentally demonstrated 5-bit multilevel programming with a high cyclability $>$1600 switching events (Fig. 5.c) [109]. Besides, all-optical SNNs using chalcogenide PCM GeSbTe (GST) also have been reported [32], and more details will be discussed in Section. III-C2. A brief summary of modulation techniques with Si waveguide is shown in Table I.
Figure 5: Modulation techniques of photonic-electronic devices on PICs. (a) Optical micrograph and the normalized transmission curve of a thermo-optic MZI. Each arm of the MZI incorporates a microheater as a phase shifter [89]. While both arms can be used to modulate the output signal, typically, one is used to set the modulator output at the quadrature point for a high extinction ratio in monotonic modulation. (b) Schematics and micrographs of free-carrier effect-based modulators working in carrier-injection mode [110], carrier-depletion mode [95], and by the MOSCAP-driven tuning mechanism [98]. (c) Top, optical micrograph of an MRR modulator loaded with 10 $\mu$m long 20-nm-thick Sb2S3 and doped silicon PIN heater. Bottom, normalized transmission spectra of MRR when switching between two phases of Sb2S3 [109].
@@
Thermo-optics $\sim$0.023 $<$1 $\sim$4 KHz $\sim$1.36 mW* $\sim$170 $\mu$m MZM [89]
Carrier Injection 0.036 12 16 GHz 10 Gb/s 51 mW* / 5 pJ/bit 200 $\mu$m MZM [92]
Carrier Depletion 1.62–2.05 3.9 30 GHz 44 Gb/s 2.84 pJ/bit 1000 $\mu$m MZM [95]
Si-MOSCAP $\sim$0.8 9 50 GHz $>$100 Gb/s R = 15 $\mu$m MRR [96]
III-V-MOSCAP 0.09 1 2.2 GHz 32 Gb/s 250 $\mu$m MZM [98]
Graphene-MOSCAP 0.28 $\sim$7 5 GHz 10 Gb/s 1 pJ/bit 300 $\mu$m MZM [111]
@@
III-B Implementations of Photonic Tensor Cores
As illustrated in Fig. 4, the interconnections in ANNs can be conceptualized as weights akin to synaptic coupling coefficients in biological systems. This analogy extends to representing these connections through tensor operations, thereby abstracting the complex interactions in a computationally manageable form. Building upon the aforementioned modulation mechanisms, various active devices and encoding mechanisms have been utilized effectively in photonic neurons. This section categorizes ONNs from multiple perspectives, which aims to provide a comprehensive analysis and comparison of the implementations of photonic tensor cores (PTCs) from diverse optical components to the architectural design.
@@ III-B1 Coherent vs incoherent ONN
From the viewpoint of signal properties, ONNs can be classified into coherent and incoherent systems. Within coherent ONNs, both weights and inputs can be encoded in the complex plane, allowing multiplication through lossless interference. For instance, a pair of beam splitters and phase shifters in the form of MZIs are widely adopted for conducting linear operations in coherent ONNs due to the programmable amplitude and phase response [112]. The schematic of a 2$\times$2 MZI is shown in Fig. 6.a. Assuming that each beam splitter is an ideal 50:50 directional coupler (i.e., with a coupling coefficient $\kappa=1/\sqrt{2}$), the transition matrix of the MZI can be represented as:
@@
Here, $\phi_{1}$ and $\phi_{2}$ represent the phase changes as signals traverse two arms. In some cases, $\phi_{1}$ and $\phi_{2}$ are denoted as $\Delta\phi=\phi_{1}-\phi_{2}$ and 0, respectively, given that the response is predominantly governed by the phase difference between its two arms. If only one input port receives a signal $\bm{E}_{\textbf{in}}$, the field and power intensities at the output port can be expressed as follows:
@@
This implies that the phase and amplitude of output can be modulated by varying $\Delta\phi$ from 0 to $\pi$. The smooth transmission curve, stemming from the cosine term, enables high-resolution analog computing and enhances noise resistance. In a notable case, the two arms are encoded by opposite phase shifts, represented as $\phi\pm\Delta\phi/2$ for $\phi_{1}$ and $\phi_{2}$, respectively. When incorporated into Eq. (4), the Mach-Zehnder modulator (MZM) working in “push-pull” mode allows for the independent modulation of field intensity while maintaining a constant phase.
@@
$U(2)$ $=\overbrace{\left[\begin{array}[]{cc}e^{-j\theta}&0\\[4.0pt] 0&1\\ \end{array}\right]}^{\text{Phase Shifter}}\cdot\overbrace{\frac{1}{2}\left[\begin{array}[]{cc}e^{-j\Delta\phi}-1&j(e^{-j\Delta\phi}+1)\\[6.0pt] j(e^{-j\Delta\phi}+1)&-(e^{-j\Delta\phi}-1)\\ \end{array}\right]}^{\text{MZI}}$ (11)
@@
Unitary matrices of rank $N$ can be decomposed into sets of $U(2)$ blocks, which can be implemented using cascaded 2$\times$2 MZIs to form a mesh structure. Despite the unitary nature of the transition matrix in an MZI-mesh, an arbitrary weight matrix for an ONN can be implemented by singular vector decomposition (SVD). The SVD approach was thoroughly discussed by Miller et al. in 2003 and first experimentally implemented in ONNs by Shen et al. as shown in Fig. 6.b [113, 28]. Specifically, an arbitrary real-valued matrix $\bm{M}$ can be decomposed into the form $\bm{U\Sigma V^{\dagger}}$, where $\bm{\Sigma}$ is a diagonal matrix and the remaining two are unitary matrices, which can be implemented with a set of tunable attenuators and an MZI mesh, respectively.
Another characteristic of the devices based on phase modulation is their broad spectral bandwidth. During the modulation process, the phase change $\psi$ can be expressed as follows:
@@
Thanks to the phase encoding mechanism, coherent ONNs can perform multiplication with full-range input operands. Besides the SVD approach, a more intuitive example is that the sign and absolute value of the scalar can be encoded by the phase shifter and MZM separately. Following this methodology, Ref [115] presents a photonic neuron architecture enabling full-range dot product operations. However, this architecture also faces limitations in terms of complexity and scalability. Specifically, implementing SVD required a matrix pre-processing step for phase mapping, which consumes extra time and power. In addition, the number of MZIs escalates quadratically with matrix size, leading to scalability issues due to the larger footprint and accumulated losses. Moreover, the phase control of the coherence network presents a challenge as well. Within the MZI mesh, each $U(2)$ operation demands a minimum of two active phase modulators. Without a reference signal, phase calibration becomes more complex than amplitude calibration (which can be straightforwardly measured using photodetectors), especially considering the fabrication deviation across individual devices and waveguides. However, some studies demonstrate the feasibility of addressing the phase calibration issue through on-chip training [116, 117].
Incoherent PICs offer an alternative approach to execute tensor operations. In the absence of coherence requirements, an incoherent ONN enables broader flexibility in employing multiplexing techniques. Because of its wavelength-dependent transmission characteristics, MRRs are widely employed for encoding weights and inputs in WDM-based incoherent ONNs. From the perspective of physical design, the small footprint of MRR (achieving a radius <10 µm) affords a compact layout and an enhanced scalability. As shown in Fig. 6.c, the add-drop MRR is a four-port optical device that consists of a microring evanescently coupled to two bus waveguides. The transfer characteristics at the through port and drop port can be expressed as Eq. (16) and Eq. (17).